OpenStudy (falling_in_katt):

The lengths of the sides of a triangle are 3, 3, . Can the triangle be a right triangle? YES or NO

Mehek (mehek14):

what's the last length?

OpenStudy (anonymous):

Yee

OpenStudy (falling_in_katt):

3\/2

OpenStudy (mathstudent55):



OpenStudy (anonymous):

\[\large 3\sqrt{2}\] right?

OpenStudy (falling_in_katt):

yes

OpenStudy (anonymous):

Yes, this fits under that of a 45-45-90 triangle. The two legs have equal lengths and are both 45 degrees. The hypotenuse will be \(Leg~Length\sqrt{2}\) and since the legs are 3 and the hypotenuse is \(3\sqrt{2}\) it is a right triangle.

OpenStudy (falling_in_katt):

Thank you!

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]